OpenStudy (anonymous):

cos(x/5)sin(x/5)=1/2[sin(2x/5)] true or false... i think its true but im not sure plz help if u an thANKS :)

OpenStudy (lgbasallote):

hmmm imagine x/5 = A so... cos(A)sin(A) = 1/2 sin(2A) multiply both sides by 2 (to get rid of 1/2 on the right side) 2cos(A)sin(A) = sin(2A) is it? :D

OpenStudy (anonymous):

thanks :)